OpenStudy (anonymous):

(-10) to the fourth power (13)

OpenStudy (anonymous):

umm 400

OpenStudy (saifoo.khan):

\[(-10)^4 \times 13\]

OpenStudy (saifoo.khan):

that?

OpenStudy (anonymous):

no~

OpenStudy (anonymous):

-32

OpenStudy (saifoo.khan):

o_O

OpenStudy (anonymous):

o_O

OpenStudy (saifoo.khan):

O_o

OpenStudy (anonymous):

Answer =D

OpenStudy (saifoo.khan):

HOWWWWWWWWW? :O

OpenStudy (anonymous):

I dont know !!! MY BRAIN HURTS NOW D:

OpenStudy (saifoo.khan):

DAng, get some music.

OpenStudy (anonymous):

=D soo hehee whats the answer :P -GIVES GUMMY WORM - :D

OpenStudy (saifoo.khan):

lol.

OpenStudy (saifoo.khan):

\[(-10)^4(13) = 1.3 \times 10^{14}\]

OpenStudy (anonymous):

Please!!

OpenStudy (saifoo.khan):

that's the answer!!! :O

OpenStudy (anonymous):

no its a number :O FOR THE ANSWER

OpenStudy (saifoo.khan):

Im lost! :/

OpenStudy (saifoo.khan):

\[(-10)^4=-10,000\]

OpenStudy (anonymous):

No!! ok so whats is (-10)to the 4th power(13)=?

OpenStudy (anonymous):

=D

OpenStudy (saifoo.khan):

\[(-10)^4(13) = 1.3 \times 10^{14}\]

OpenStudy (saifoo.khan):

are u lost as well? LOL.

OpenStudy (anonymous):

Yees!!

OpenStudy (anonymous):

omg im soo confused

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

haha, where are u stuck?

OpenStudy (anonymous):

Well lie whats the answer

OpenStudy (anonymous):

like

OpenStudy (saifoo.khan):

lol, im stuck that wht's the question.

OpenStudy (anonymous):

lol (-10) to the fourth power (13)=

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

130,000

OpenStudy (saifoo.khan):